OpenStudy (anonymous):

80% of a number is 96 What is 25% of that same number? A.20 B.30 C.40 D.120 -will give medal @aureyliant @baconbaby1334 @dianolove @RosieF

OpenStudy (anonymous):

i think you should divide 96 by 0.80

OpenStudy (anonymous):

ok its 90 divde bye 0.80 cause %80 and the number is 90 so yea

OpenStudy (anonymous):

ok let me try and see

OpenStudy (anonymous):

the answer is b.30 @Angelsoft

OpenStudy (anonymous):

80/100 = 96/x 80x=9600 8x=960 x=120 120/4 30

OpenStudy (anonymous):

up u right i was thinking liuke somthing is worng here it canit be D i know its B but i forgot but u r u its B

OpenStudy (anonymous):

thx every one or ur help!

OpenStudy (anonymous):

since we found out what number the question was talking about (120) i find 25% of 120, which is 120*0.25, which is 30 that's how i found out

OpenStudy (anonymous):

who agrees when u ask a question and theres a lot of ppl saying the right answer but u have to give out 1 medal.it stinks I wish I can give it to all of yall!

OpenStudy (anonymous):

one solution is i give a medal to someone, then that other someone gives a medal to someone else

OpenStudy (anonymous):

i gave @Rubyjadeluv

OpenStudy (anonymous):

a medal

OpenStudy (anonymous):

ok I gave u one then :) thx for ur help @itsthatguy!

OpenStudy (anonymous):

so @Rubyjadeluv should give a medal to @dianolove so we have medals

OpenStudy (anonymous):

anyway, i have to log out

OpenStudy (anonymous):

goo idea thank so much add me if u want!